# All Languages Matter: Evaluating LMMs on Culturally Diverse 100 Languages

### v2

@@ Abstract
Existing Large Multimodal Models (LMMs) generally focus on only a few regions and languages. As LMMs continue to improve, it is increasingly important to ensure they understand cultural contexts, respect local sensitivities, and support low-resource languages, all while effectively integrating corresponding visual cues. In pursuit of culturally diverse global multimodal models, our proposed All Languages Matter Benchmark (ALM-bench) represents the largest and most comprehensive effort to date for evaluating LMMs across 100 languages. ALM-bench challenges existing models by testing their ability to understand and reason about culturally diverse images paired with text in various languages, including many low-resource languages traditionally underrepresented in LMM research. The benchmark offers a robust and nuanced evaluation framework featuring various question formats, including true/false, multiple choice, and open-ended questions, which are further divided into short and long-answer categories. ALM-bench design ensures a comprehensive assessment of a model’s ability to handle varied levels of difficulty in visual and linguistic reasoning. To capture the rich tapestry of global cultures, ALM-bench carefully curates content from 13 distinct cultural aspects, ranging from traditions and rituals to famous personalities and celebrations. Through this, ALM-bench not only provides a rigorous testing ground for state-of-the-art open and closed-source LMMs but also highlights the importance of cultural and linguistic inclusivity, encouraging the development of models that can serve diverse global populations effectively. Our benchmark is publicly available.
@@ 1 Introduction
In recent years, large multimodal models (LMMs) have seen significant progress across a variety of vision-language tasks [2, 41, 5, 10, 28, 36, 30]. Despite these advances, a major limitation remains in the models’ capacity to understand and respond accurately to cultural and linguistic diversity [37]. Current LMMs perform well in widely spoken languages with substantial training data but struggle with low-resource languages and nuanced cultural contexts, limiting their effectiveness on a global scale [29].
[…]
Our contributions can be summarized as follows:
2 Related Work
To align large language models (LLMs) with human preferences, various approaches have been developed [50, 35, 39]. Over time, these methods are extended to enable LLMs to be sensitive to geo-diverse cultural values [1, 24]. This evolution has broadened their applicability to both multilingual and multimodal LLMs, enabling greater engagement with users from diverse linguistic and cultural backgrounds. Despite this progress, LMM benchmarks often fail to adequately address cultural aspects, especially for low-resource languages. Models such as GPT-4V [2] and Gemini 1.5 Pro [41] performs well on existing multimodal benchmarks, but these evaluations often lack in capturing the intricacies of multilingual and culturally diverse interactions. Open-source models, in particular, face challenges in accurately representing geodiverse values (e.g., local beliefs, rituals, and traditions), underscoring the need for a culturally aware evaluation framework for LMMs (see Tab. 1).
@@ 3 ALM Benchmark
ALM-bench provides global coverage, featuring both generic and culture-specific image samples that preserve the nuances of native languages in VQAs (Fig. 4). It includes diverse QA formats such as MCQs, True/False, and open-ended VQA (long and short) under 19 domains, grouped into generic and cultural categories Tab. A.1 (suppl. material).
@@ 3.1 Data Collection and Annotation
We note several challenges in GPT-4o’s translations across 100 languages. Specifically, GPT-4o demonstrates difficulty in generating culturally grounded QA pairs across diverse languages. To analyze these issues, mistakes from the GPT-4o model in translations are categorized into four types: semantic error, cultural error, language error, and grammatical error. We sample 57 question-answer pairs from 51 randomly selected languages and plotted error distribution in Fig. 5. Notably, GPT-4o encounters more issues with semantic and grammatical accuracy when translating into different languages. We show examples of such errors in Fig. A.2 (suppl. material).
@@
Figure 6: ALM-bench Performance comparison of different open and closed-sourced models (y-axis) on the 100 languages (x-axis) of our ALM-bench. The performance is represented as an average accuracy across all questions in a language. The actual performance of a model on a language is shown in each respective box, where the higher accuracy is highlighted with a high color intensity. The closed-source propriety models generally perform better across languages compared to their open-sourced counterparts. The performance on several high resource languages (e.g., English, French, Chinese and Spanish) is generally higher throughout different models, whereas all open-source models struggle on low resource languages (e.g., Amharic, Kinyarwanda, Burmese) Overall, GPT-4o and GLM-4V-9B performs better in terms of closed-source and open-source models, respectively. Best viewed zoomed in.
@@ 4 Benchmarking LMMs on ALM-bench
As discussed earlier, ALM-bench comprises four different question types: MCQs, T/F, Short and Long VQA. Therefore, we employ different prompts for each type of question. We use GPT-4o as a judge and prompt it to score responses of various question types with different criterion. We use accuracy for MCQs and T/F, correctness for short VQAs (SVQAs), and consistency, fluency, and relevance for long VQAs (LVQAs).
Here, correctness refers to how closely the model’s output aligns with the ground-truth [26]. For LVQAs, the consistency assesses whether the predicted answer is coherent across the entire generated output [20]. The fluency metric measures the naturalness and readability of model prediction [38, 42], whereas the relevancy metric determines whether model prediction provides answers directly related to ground truth [7]. To ensure fair evaluation for decision-making questions (T/F and MCQs) using GPT-4o as a judge, we also show consistent performance using LLaMA-3.8.1B-Instruct [14] in Tab. A.2 (suppl. material).
Overall Results: Fig. 6 presents the per-language performance of 16 recent LMMs on the ALM-bench. The results offer a number of insights: (a) the closed-source propriety models (GPT-4o [2] and Gemini-1.5-Pro [41]) generally perform better across the 100 languages, compared to their open-source counterparts. The best closed-source model, GPT-4o, achieves an overall accuracy of 78.8% compared to 51.9% overall accuracy obtained by the best performing open-source model GLM-4V-9B [16]. Both closed-source models struggle with several low-resource languages (e.g., Amharic, Kinyarwanda, Burmese, and Sanskrit). For instance, the performance of GPT-4o significantly drops from 88.4% for the English language to 50.8% for the Amharic language. (b) Several open-source models (GLM-4V-9B [16], Qwen2-VL [44], Molmo [13], LLaVA-OneVision [23]) achieve comparable overall performance.
Similar to their closed-source counterparts, open-source models also struggle with low-resource languages. For instance, the performance of GLM-4V-9B [16] dramatically drops from 80.3% for the English language to 15.6% for the Amharic language.
Fig. 2a presented earlier further shows the overall performance breakdown in terms of high-and low-resource languages across different models. It is worth mentioning that our ALM-bench comprises 50 high-resource and 50 low-resource languages following the definition of low-resource languages as in [11]. (c) Fig. 2a reveals a substantial performance gap between high-resource and low-resource languages consistent across different models, with GPT-4o showing a 6% drop. This performance gap extends to more than 8% in case of some open-source models (e.g., GLM-4V-9B [16], Qwen2-VL [44]). A notable exception is the performance of closed-source Gemini-1.5-Pro [41], which does not deteriorate significantly for low-resource languages.
Next, we present further results analysis.
@@
Models With Country Info. Without Country Info.
GPT-4o 83.57% 80.96%
Gemini-1.5-Pro 81.52% 76.19%
[…]
Qwen2-VL 53.97% 52.57%
Table 2: Performance of LMMs with and without additional country location information. Proprietary models show a notable performance boost of 2.6% to 5% when location-aware prompts are used, while open-source models exhibit a marginal improvement.
@@
Comparison across language families: We also analyze the performance of LMMs by grouping the different languages with respect to 15 families taken from Britannica [^1] https://www.britannica.com/browse/Languages as shown in Fig. 9.
Results show that performance on several African (Atlantic-Congo) languages such as Igbo, Kinyarwanda, Shona, Swahili, and Yoruba is inferior compared to several Asian (e.g., Chinese, Korean, Vietnamese) and Western languages (e.g., English, French, German).
@@
Are LMMs Culturally Aware? We study the cultural understanding of LMMs on our 13 domains and report results in Fig. 2b. Overall, GPT-4o achieves the best results with an overall score of 80.3%. However, we observe a significant variation in performance across different cultural domains. For instance, GPT-4o achieves 83.7% on Education and heritage but drops to 72.7% on Notable key figures category. We conjecture this variation in performance likely arises since the higher-performing categories are covered in LMM training datasets, leading to better representation learning. In contrast, categories like Notable Key Figures and Customs are often culturally specific and nuanced across languages and regions.
Impact of location-aware information in prompts: We show the performance of LMMs when evaluated with additional country-specific information in Tab. 2. We randomly select 50 languages from the ALM-bench and include country information in the prompts to assess any potential performance gains. Here, we observe that closed-source models, including GPT-4o and Gemini-1.5-Pro, better utilize the added geographic context and indicate higher cultural specificity across languages than the open-source models.
@@ 7 Acknowledgment
We would like to thank Tafar Mab for providing high-quality annotations for Afrikaans and Kinyarwanda languages, Georgios Ioannides for verifying Greek language, and Feno Heriniaina Rabevohitra for annotating Malagasy language. We also thank Yuhao Li for reviewing and assisting with the annotations for Chinese Simplified language. Finally, we extend our gratitude to the anonymous reviewers for their invaluable suggestions and feedback, which helped improve this paper.
@@ Appendix A ALM-Bench Categories
Our ALM-bench benchmark dataset consists of 19 categories. Among these categories, we include Food, Lifestyle, Religion, Architecture, Sports, Customs and Heritage from [37, 31] with some minor modifications to it. We further include six additional categories including Literature, featuring prominent authors, poets, and their notable works; Music, showcasing traditional music and dance through visual samples; Festivals, covering major cultural celebrations; Economy, representing local industries and businesses; Media, highlighting cultural icons, entertainment figures, and popular TV shows; and Notable Key Figures, featuring influential historical leaders who serve as representatives for country-language association. Table. A. A.1 entails all our ALM-bench categories.
Following [3], we adopt an approach to group cultural attributes based on the country for each language. Additionally, we group culturally grounded elements representing shared knowledge, values, and objectives among the people in that culture that are collectively understood. We curated a culturally diverse dataset by collecting data from many cultural aspects (see Fig. A.1) and manually curated the caption for each image from web-sources. We also show the most frequent words from these categories in Fig. A.15.
@@ 19. Notable Key Figures
Table A.1: The 19 categories present in our ALM-bench dataset. Note that Food Items is sourced from generic categories while the Food subset is sourced from the cultural category from our ALM-bench.
@@ Appendix B Annotators Demographics
Each annotator was given detailed instructions along with examples to guide them in curating high-quality datasets. We also hosted a UI interface as well as gave them the option to directly use the dataset. An example of our annotation platform is shown in Fig. A.4. Annotators who made substantial contributions are recognized as co-authors of this paper to support high-quality language verification.
We summarize the annotator demographics in Fig. A.3. To summarize, we show the Age and Gender distribution of our volunteers (top-left) indicating the presence of one-fourth of Female annotators and over 46% fall in the 18-25 age limit. We show their levels of language proficiency (top-right). Notably, over 80% of the annotators were either native speakers or bilingual, with an additional 14.8% at a proficient level in their chosen language. This high level of expertise ensures that our cultural samples have been meticulously reviewed, filtered, and refined by knowledgeable volunteers. Additionally, we pre-select annotators who are accustomed to the cultural values of their languages. Moreover, (bottom-left) depicts that over 88.5% are culturally familiar with their languages, and bottom-right highlights that 87.9% of the annotators have lived in the country where we have crafted the cultural samples, making them ideal participants for participating in this research effort.
@@ Appendix C Guidelines for Data Verification
Step 1 (Question Relevance): Verify that the question is phrased such that it cannot be answered without the use of a Vision-Language Model (VLM) analyzing the image. For example, a question like “Where is Aura Church located?” can be answered without any visual input. A more appropriate question would be, “Where is the church shown in the image located?
@@ Appendix G Instruction Prompt for QA Generation
Fig. A.5 illustrates the prompt messages used to guide GPT-4o in generating culturally relevant QA pairs. Each prompt includes an image, its manually added caption, and the search query used during image scraping. The model is instructed to generate two short QA pairs, two multiple-choice QA pairs, one true/false question, and one long QA pair. Detailed instructions ensure that the generated QAs are culturally grounded and specific to the image, avoiding answers based solely on common knowledge.
@@ Appendix H Instruction Prompt for QA Translations
After the curation of question-answer pairs using GPT-4o, we use the same open-source model to translate the English QA pairs into the native languages. We show the prompt that we used for this task in Fig. A.6.
@@ Appendix I Instruction Prompt for LMM Answer Generation
We conduct a comprehensive study by evaluating various state-of-the-art Large Multimodal Models (LMMs), including both open-weight and closed-weight models, on our ALM-bench benchmark. We used different prompts for each question type as shown in Fig. A.7. We prompt all questions with English system instruction as highlighted by [19] that the use of prompts in English results in the best performance. Finally, we score each model’s generated answer with the human-annotated ground truth answers through a scoring system to assess each model’s performance on the benchmark.
@@
In addition to using GPT-4o as a judge in scoring the predicted answers, we also do scoring with multiple other LMMs to ensure a fair evaluation. Results are reported in Table. A. A.2
@@
Next, we present some qualitative examples showing various question types for each image sample from our dataset and in Figs. A.9, A.10, A.11, A.11, A.13, and A.14.

### v3

@@ Abstract
11footnotetext: corresponding author: ashmal.vayani@ucf.edu
Existing Large Multimodal Models (LMMs) generally focus on only a few regions and languages. As LMMs continue to improve, it is increasingly important to ensure they understand cultural contexts, respect local sensitivities, and support low-resource languages, all while effectively integrating corresponding visual cues. In pursuit of culturally diverse global multimodal models, our proposed All Languages Matter Benchmark (ALM-bench) represents the largest and most comprehensive effort to date for evaluating LMMs across 100 languages. ALM-bench challenges existing models by testing their ability to understand and reason about culturally diverse images paired with text in various languages, including many low-resource languages traditionally underrepresented in LMM research. The benchmark offers a robust and nuanced evaluation framework featuring various question formats, including true/false, multiple choice, and open-ended questions, which are further divided into short and long-answer categories. ALM-bench design ensures a comprehensive assessment of a model’s ability to handle varied levels of difficulty in visual and linguistic reasoning. To capture the rich tapestry of global cultures, ALM-bench carefully curates content from 13 distinct cultural aspects, ranging from traditions and rituals to famous personalities and celebrations. Through this, ALM-bench not only provides a rigorous testing ground for state-of-the-art open and closed-source LMMs but also highlights the importance of cultural and linguistic inclusivity, encouraging the development of models that can serve diverse global populations effectively. Our benchmark is publicly available at https://mbzuai-oryx.github.io/ALM-Bench/.
@@ 1 Introduction
In recent years, large multimodal models (LMMs) have seen significant progress across a variety of vision-language tasks [2, 42, 5, 9, 28, 36]. Despite these advances, a major limitation remains in the models’ capacity to understand and respond accurately to cultural and linguistic diversity [37]. Current LMMs perform well in widely spoken languages with substantial training data but struggle with low-resource languages and nuanced cultural contexts, limiting their effectiveness on a global scale [29].
[…]
Our contributions can be summarized as follows:
2 Related Work
To align large language models (LLMs) with human preferences, various approaches have been developed [50, 35, 39]. Over time, these methods are extended to enable LLMs to be sensitive to geo-diverse cultural values [1, 24, 31, 40]. This evolution has broadened their applicability to both multilingual and multimodal LLMs, enabling greater engagement with users from diverse linguistic and cultural backgrounds. Despite this progress, LMM benchmarks often fail to adequately address cultural aspects, especially for low-resource languages. Models such as GPT-4V [2] and Gemini 1.5 Pro [42] performs well on existing multimodal benchmarks, but these evaluations often lack in capturing the intricacies of multilingual and culturally diverse interactions. Open-source models, in particular, face challenges in accurately representing geodiverse values (e.g., local beliefs, rituals, and traditions), underscoring the need for a culturally aware evaluation framework for LMMs (see Tab. 1).
@@ 3 ALM Benchmark
ALM-bench provides global coverage, featuring both generic and culture-specific image samples that preserve the nuances of native languages in VQAs (Fig. 4). It includes diverse QA formats such as MCQs, True/False, and open-ended VQA (long and short) under 19 domains, grouped into generic and cultural categories Tab. A. 1. (suppl. material).
@@ 3.1 Data Collection and Annotation
We note several challenges in GPT-4o’s translations across 100 languages. Specifically, GPT-4o demonstrates difficulty in generating culturally grounded QA pairs across diverse languages. To analyze these issues, mistakes from the GPT-4o model in translations are categorized into four types: semantic error, cultural error, language error, and grammatical error. We sample 57 question-answer pairs from 51 randomly selected languages and plotted error distribution in Fig. 5. Notably, GPT-4o encounters more issues with semantic and grammatical accuracy when translating into different languages. We show examples of such errors in Fig. A. 2. (suppl. material).
@@
Figure 6: ALM-bench Performance comparison of different open and closed-sourced models (y-axis) on the 100 languages (x-axis) of our ALM-bench. The performance is represented as an average accuracy across all questions in a language. The actual performance of a model on a language is shown in each respective box, where the higher accuracy is highlighted with a high color intensity. The closed-source propriety models generally perform better across languages compared to their open-sourced counterparts. The performance on several high-resource languages (e.g., English, French, Chinese, and Spanish) is generally higher throughout different models, whereas all open-source models struggle on low resource languages (e.g., Amharic, Kinyarwanda, Burmese) fs QAW a Overall, GPT-4o and GLM-4V-9B performs better in terms of closed-source and open-source models, respectively. Best viewed zoomed in.
@@ 4 Benchmarking LMMs on ALM-bench
As discussed earlier, ALM-bench comprises four different question types: MCQs, T/F, Short and Long VQA. Therefore, we employ different prompts for each type of question. We use GPT-4o as a judge and prompt it to score responses of various question types with different criterion. We use accuracy for MCQs and T/F, correctness for short VQAs (SVQAs), and consistency, fluency, and relevance for long VQAs (LVQAs). Here, correctness refers to how closely the model’s output aligns with the ground-truth [26]. For LVQAs, the consistency assesses whether the predicted answer is coherent across the entire generated output [19]. The fluency metric measures the naturalness and readability of model prediction [38], whereas the relevancy metric determines whether model prediction provides answers directly related to ground truth [6]. To ensure fair evaluation for decision-making questions (T/F and MCQs) using GPT-4o as a judge, we also show consistent performance using LLaMA-3.8.1B-Instruct [13] in Tab. A.2. (suppl. material).
Overall Results: Fig. 6 presents the per-language performance of 16 recent LMMs on the ALM-bench. The results offer a number of insights: (a) the closed-source propriety models (GPT-4o [2] and Gemini-1.5-Pro [42]) generally perform better across the 100 languages, compared to their open-source counterparts. The best closed-source model, GPT-4o, achieves an overall accuracy of 78.8% compared to 51.9% overall accuracy obtained by the best performing open-source model GLM-4V-9B [15]. Both closed-source models struggle with several low-resource languages (e.g., Amharic, Kinyarwanda, Burmese, and Sanskrit). For instance, the performance of GPT-4o significantly drops from 88.4% for the English language to 50.8% for the Amharic language. (b) Several open-source models (GLM-4V-9B [15], Qwen2-VL [44], Molmo [12], LLaVA-OneVision [23]) achieve comparable overall performance. Similar to their closed-source counterparts, open-source models also struggle with low-resource languages. For instance, the performance of GLM-4V-9B [15] dramatically drops from 80.3% for the English language to 15.6% for the Amharic language. Fig. 2(a) presented earlier further shows the overall performance breakdown in terms of high-and low-resource languages across different models. It is worth mentioning that our ALM-bench comprises 50 high-resource and 50 low-resource languages following the definition of low-resource languages as in [10]. (c) Fig. 2(a) reveals a substantial performance gap between high-resource and low-resource languages consistent across different models, with GPT-4o showing a 6% drop. This performance gap extends to more than 8% in case of some open-source models (e.g., GLM-4V-9B [15], Qwen2-VL [44]). A notable exception is the performance of closed-source Gemini-1.5-Pro [42], which does not deteriorate significantly for low-resource languages. Next, we present further results analysis.
@@
Comparison across language families: We also analyze the performance of LMMs by grouping the different languages with respect to 15 families taken from Britannica***https://www.britannica.com/browse/Languages as shown in Fig. 9. Results show that performance on several African (Atlantic-Congo) languages such as Igbo, Kinyarwanda, Shona, Swahili, and Yoruba is inferior compared to several Asian (e.g., Chinese, Korean, Vietnamese) and Western languages (e.g., English, French, German).
@@
Are LMMs Culturally Aware? We study the cultural understanding of LMMs on our 13 domains and report results in Fig. 2(b). Overall, GPT-4o achieves the best results with an overall score of 80.3%. However, we observe a significant variation in performance across different cultural domains. For instance, GPT-4o achieves 83.7% on Education and heritage but drops to 72.7% on Notable key figures category. We conjecture this variation in performance likely arises since the higher-performing categories are covered in LMM training datasets, leading to better representation learning. In contrast, categories like Notable Key Figures and Customs are often culturally specific and nuanced across languages and regions.
Impact of location-aware information in prompts: We show the performance of LMMs when evaluated with additional country-specific information in Tab. LABEL:tab:country_info_performance. We randomly select 50 languages from the ALM-bench and include country information in the prompts to assess any potential performance gains. Here, we observe that closed-source models, including GPT-4o and Gemini-1.5-Pro, better utilize the added geographic context and indicate higher cultural specificity across languages than the open-source models.
@@ 7 Acknowledgment
We would like to thank Tafar Mab for providing high-quality annotations for Afrikaans and Kinyarwanda languages, Georgios Ioannides for verifying Greek language, Marek Suppa for annotating and verifying Czech, Slovak and Polish language, and Feno Heriniaina Rabevohitra for annotating Malagasy language. We also thank Yuhao Li for reviewing and assisting with the annotations for Chinese Simplified language. Finally, we extend our gratitude to the anonymous reviewers for their invaluable suggestions and feedback, which helped improve this paper. The computations were enabled by resources provided by NAISS at Alvis partially funded by Swedish Research Council through grant agreement no. 2022-06725, LUMI hosted by CSC (Finland) and LUMI consortium, and by Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the NSC. This work was partially supported by by the Swedish Research Council (2022-04266) and from KAW (DarkTree project; 2024.0076).
@@ Appendix A ALM-Bench Categories
Our ALM-bench benchmark dataset consists of 19 categories. Among these categories, we include Food, Lifestyle, Religion, Architecture, Sports, Customs and Heritage from [37, 30] with some minor modifications to it. We further include six additional categories including Literature, featuring prominent authors, poets, and their notable works; Music, showcasing traditional music and dance through visual samples; Festivals, covering major cultural celebrations; Economy, representing local industries and businesses; Media, highlighting cultural icons, entertainment figures, and popular TV shows; and Notable Key Figures, featuring influential historical leaders who serve as representatives for country-language association. Table A. A.1 entails all our ALM-bench categories.
Following [3], we adopt an approach to group cultural attributes based on the country for each language. Additionally, we group culturally grounded elements representing shared knowledge, values, and objectives among the people in that culture that are collectively understood. We curated a culturally diverse dataset by collecting data from many cultural aspects (see Fig. A. A.1) and manually curated the caption for each image from web-sources. We also show the most frequent words from these categories in Fig. A. A.15.
@@ 19. Notable Key Figures
Table A.1: The 19 categories present in our ALM-bench dataset. Note that Food Items and Food appear in both generic and cultural categories, but have a different context as described in Sec. F.
@@ Appendix B Annotators Demographics
Each annotator was given detailed instructions along with examples to guide them in curating high-quality datasets. We also hosted a UI interface as well as gave them the option to directly use the dataset. An example of our annotation platform is shown in Fig. A. A.4. Annotators who made substantial contributions are recognized as co-authors of this paper to support high-quality language verification.
We summarize the annotator demographics in Fig. A. A.3. To summarize, we show the Age and Gender distribution of our volunteers (top-left) indicating the presence of one-fourth of Female annotators and over 46% fall in the 18-25 age limit. We show their levels of language proficiency (top-right). Notably, over 80% of the annotators were either native speakers or bilingual, with an additional 14.8% at a proficient level in their chosen language. This high level of expertise ensures that our cultural samples have been meticulously reviewed, filtered, and refined by knowledgeable volunteers. Additionally, we pre-select annotators who are accustomed to the cultural values of their languages. Moreover, (bottom-left) depicts that over 88.5% are culturally familiar with their languages, and bottom-right highlights that 87.9% of the annotators have lived in the country where we have crafted the cultural samples, making them ideal participants for participating in this research effort.
@@ Appendix C Guidelines for Data Verification
Step 1 (Question Relevance): Verify that the question is phrased such that it cannot be answered without the use of a Vision-Language Model (VLM) analyzing the image. For example, a question like “Where is Aura Church located?” can be answered without any visual input. A more appropriate question would be, “Where is the church shown in the image located?"
@@ Appendix G Instruction Prompt for QA Generation
Fig. A. A.5 illustrates the prompt messages used to guide GPT-4o in generating culturally relevant QA pairs. Each prompt includes an image, its manually added caption, and the search query used during image scraping. The model is instructed to generate two short QA pairs, two multiple-choice QA pairs, one true/false question, and one long QA pair. Detailed instructions ensure that the generated QAs are culturally grounded and specific to the image, avoiding answers based solely on common knowledge.
@@ Appendix H Instruction Prompt for QA Translations
After the curation of question-answer pairs using GPT-4o, we use the same open-source model to translate the English QA pairs into the native languages. We show the prompt that we used for this task in Fig. A. A.6.
@@ Appendix I Instruction Prompt for LMM Answer Generation
We conduct a comprehensive study by evaluating various state-of-the-art Large Multimodal Models (LMMs), including both open-weight and closed-weight models, on our ALM-bench benchmark. We used different prompts for each question type as shown in Fig. A. A.7. We prompt all questions with English system instruction as highlighted by [18], suggesting that the use of prompts in English results in the best performance. Finally, we score each model’s generated answer with the human-annotated ground truth answers through a scoring system to assess each model’s performance on the benchmark.
@@
In addition to using GPT-4o as a judge in scoring the predicted answers, we also do scoring with multiple other LMMs to ensure a fair evaluation. Results are reported in Table A. A.2.
@@
Next, we present some qualitative examples showing various question types for each image sample from our dataset and in Fig. A. A.9, A. A.10, A. A.11, A. A.11, A. A.13, and A. A.14.
Appendix M Country Specific Prompts
For the country-specific experiment as discussed in Section 4, Impact of location-aware information in prompts, Tab. 2 (main paper), we show the prompt used to conduct this experiment below. In this setting, we aim to assess how incorporating location-aware context affects vision-language model outputs, particularly in terms of potential cultural biases. To do this, we construct prompts using structured triplets comprising (language, country, cultural domain) — for example, (“Hindi,” “India,” “Food”) or (“Japanese,” “Japan,” “Religion”). This allows us to systematically evaluate whether models respond differently to culturally grounded prompts, and whether location-specific context influences model behavior or introduces/reduces bias.
Location Aware Prompts prompt = ques + “Provide brief, clear responses in ‘lang’ language. The image represents the ‘cultural category’ in ‘country’.”